OpenStudy (anonymous):

Part 1: Solve the following equation using one of the methods listed below. Part 2: Using complete sentences, explain why you chose the method you used. Equation: 0 = x2 + 10x + 5 Methods: Completing the Square Factoring Quadratic Formula Graphing

OpenStudy (campbell_st):

completing the square: Half the coefficient of x 10/2 = 5 then 5^2 = 25 The perfecr square is (x+5)^2 . The adjustment for the constant term is -20 Then x^2 + 10x + 5 = (x + 5)^2 - 20 The parabola has vertex at (-5, -20) is concave up y intercept st (0,5) x - intercepts at -5 +- sr(5)